OpenStudy (anonymous):

I need help i don't how to do this can some one help me by walking me by step by step? [8.03] What are the factors of x2 + 12x + 20? (x + 4)(x + 5) (x + 2)(x + 10) (x − 4)(x − 5) (x − 2)(x − 10)

OpenStudy (texaschic101):

are you familiar with FOIL.....this is kind of reverse FOIL. You need to find 2 numbers when added together = 12, and when multiplied together = 20. Since there are positive signs, then the signs in your answer will be positive. (x + -)(x + -) Now what two numbers added together to equal 12 and multiplied together = 20 ?

OpenStudy (nurali):

x^2 + 12x + 20 x^2+10x+2x+20 x(x+10)+2(x+10) (x + 2)(x +10)

OpenStudy (texaschic101):

and there is the answer....

OpenStudy (texaschic101):

you can check your answer by using FOIL (x + 2)(x + 10) F- first terms multiplied x * x = x^2 O - outside terms multiplied x * 10 = 10x I - inside terms multiplied x * 2 = 2x L - last terms multiplied 2 * 10 = 20 now put it all together x^2 + 10x + 2x + 20 combine like terms x^2 + 12x + 20

OpenStudy (anonymous):

Oh ok thank you @texaschic101

OpenStudy (texaschic101):

YW :)

OpenStudy (anonymous):

and @Nurali

OpenStudy (nurali):

Anytime.